# The tame Deligne-Simpson problem

### v1

@@ 3.5. Variation of parabolic weights
$\mathbf{Var}_{2}:\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})\longrightarrow\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j}).$
The space $\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ parametrises parabolic Higgs bundles of type $\boldsymbol{\tau}(\underline{\nu})$ for some tuple of integers $\underline{\nu}=(\nu_{j})_{1\leq j\leq k}$; we require that the graded residue $\mathcal{O}_{j}$ at each puncture is central with $\nu_{j}+1$ distinct eigenvalues. The space $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ parametrises parabolic Higgs bundles of type $\boldsymbol{\tau}(\underline{\mu})$ for some $\underline{\mu}=(\mu_{j})_{1\leq j\leq k}$ with each $\mu_{j}\leq\nu_{j}$. We require that the following conditions are satisfied:
  • There is a degeneration of types $\sigma:\boldsymbol{\tau}(\underline{\mu})\longrightarrow\boldsymbol{\tau}(\underline{\nu})$ such that $\mathbf{d}^{\prime}=\sigma^{\ast}\mathbf{d}$.
@@
For arbitrary parabolic weights $\boldsymbol{\alpha}$, a degeneration of parabolic structure may not preserve semi-stability. The usual option is to take a generic $\boldsymbol{\alpha}$ in a small neighbourhood of $\boldsymbol{\alpha}^{\prime}$; however, generic weights do not exist if the dimension vector is divisible.
@@
Fix $e\in\mathbb{Z}$ and $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ and let $\boldsymbol{\alpha}\in\mathbb{R}^{\boldsymbol{\tau}(\underline{\nu})}$ be a parabolic weight such that
$e+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}d^{\ast}_{[j,i]}=0.$
We say that $\boldsymbol{\alpha}$ is almost generic if for any $\mathbf{d}^{\prime}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ of rank $n^{\prime}\leq n$ and any $e^{\prime}\in\mathbb{Z}$ the equality
$e^{\prime}+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}d^{\prime\ast}_{[j,i]}=0$
holds if and only if the vector $(e,(d_{[j,i]})_{[j,i]})$ is a $\mathbb{Q}$-multiple of $(e^{\prime},(d^{\prime}_{[j,i]})_{[j,i]})$ (if and only if $(e,(d^{\ast}_{[j,i]})_{[j,i]})$ is a $\mathbb{Q}$-multiple of $(e^{\prime},(d^{\prime\ast}_{[j,i]})_{[j,i]})$); we say that $\boldsymbol{\alpha}$ is generic if the above equality does not hold unless $(e^{\prime},(d^{\prime}_{[j,i]})_{[j,i]})$ is zero or is equal to $(e,(d_{[j,i]})_{[j,i]})$.
@@
Note that we allow $\alpha_{[j,i]}=\alpha_{[j,i+1]}$ in this space. Define
$I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d}):=\{\boldsymbol{\alpha}\in I^{\boldsymbol{\tau}(\underline{\nu})}_{\Delta}\mid e+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}d^{\ast}_{[j,i]}=0\},$
which is a hyperplane in $I^{\boldsymbol{\tau}(\underline{\nu})}_{\Delta}$. If a vector $(e^{\prime},\mathbf{d}^{\prime})$ is not a scalar multiple of $(e,\mathbf{d})$, then $I^{\boldsymbol{\tau}(\underline{\nu})}(e^{\prime},\mathbf{d}^{\prime})$ either does not meet $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$ or intersects with $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$ in a lower dimensional affine space, thus forming a wall in $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$. By definition, a weight $\boldsymbol{\alpha}\in I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$ is almost generic if it lies in the complement of the union of these walls.
@@
$\sigma_{\ast}:I^{\boldsymbol{\tau}(\underline{\mu})}(e,\sigma^{\ast}\mathbf{d})\longrightarrow I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d}).$
@@
Suppose that $\mathscr{E}=(E,E_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\nu})$ and that $\mathscr{E}^{\prime}=(E,E^{\prime}_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\mu})$ obtained from the former by a degeneration of parabolic structures $\sigma$. Let $\boldsymbol{\alpha}^{\prime}\in I_{\Delta}^{\boldsymbol{\tau}(\underline{\mu})}$. Then, we have
@@
Suppose that $(E,E_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\nu})$ and that $(E,E^{\prime}_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\mu})$ obtained from the former by a degeneration of parabolic structures $\sigma$. Let $\boldsymbol{\alpha}^{\prime}$ be a parabolic weight for $E^{\prime}_{[\bullet,\bullet]}$. Let $F\subset E$ be a $\Phi$-invariant vector bundle, and let $F^{\prime}_{[\bullet,\bullet]}$ and $F_{[\bullet,\bullet]}$ be the parabolic structures induced by $E^{\prime}_{[\bullet,\bullet]}$ and $E_{[\bullet,\bullet]}$ respectively. Then,
@@
    $(F,F^{\prime}_{[\bullet,\bullet]},\Phi)$ is obtained from $(F,F_{[\bullet,\bullet]},\Phi)$ by a degeneration of parabolic structures.
@@
Let $\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ and $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ be as in the beginning of this subsection, and denote their parabolic types by $\boldsymbol{\tau}(\underline{\nu})$ and $\boldsymbol{\tau}(\underline{\mu})$ respectively. We will define $\mathbf{Var}_{2}$ in two cases: ($\mathbf{Aff}$) and ($\mathbf{Aff}^{\infty}$). In the first case, we will need to choose $\boldsymbol{\alpha}$ to be almost generic. In the second case, we will choose $\boldsymbol{\alpha}$ to be generic; note that in this case, the stability condition on the Betti side of $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ is generic, and thus this space consists of stable parabolic Higgs bundles, although $\boldsymbol{\alpha}^{\prime}$ is not generic.
Lemma 3.6.
Let $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ and $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$ be the Dolbeault and Betti side of a nonabelian Hodge correspondence as defined in §3.3, §3.2 and §3.4, and let $\mathcal{M}(\mathbf{q},\mathbf{d})$ be the multiplicative quiver variety that is isomorphic to $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$. (We do not restrict ourselves to the cases ($\mathbf{Aff}$) and ($\mathbf{Aff}^{\infty}$).) Suppose that $\mathbf{d}=m\boldsymbol{\gamma}$ for some indivisible $\boldsymbol{\gamma}$ and $\mathbf{q}^{\boldsymbol{\gamma}}$ is an $l$-th primitive root of unity. Let $e\in\mathbb{Z}$ be such that
$e+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}(\sigma_{\ast}\alpha^{\prime})_{[j,i]}d^{\ast}_{[j,i]}=0.$
Then, the vector $(e,\mathbf{d})$ is indivisible if and only if $l=m$.
Proof.
Suppose that $l<m$. From the equality $\mathbf{q}^{l\boldsymbol{\gamma}}=1$ and the fact that the argument of $\xi_{[j,i]}$ is equal to $2\pi i\alpha^{\prime}_{[j,i]}$, we deduce that
(3.5.1) $\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}(\sigma_{\ast}\alpha)^{\prime}_{[j,i]}l\gamma^{\ast}_{[j,i]}$
is an interger, say $e^{\prime}$. (The integer $e$ as in the statement exists for the same reason.) Now, if $\mathbf{d}=m^{\prime}l\boldsymbol{\gamma}$, then $m^{\prime}e^{\prime}=e$, and thus $(e,\mathbf{d})$ is divisible. Conversely, suppose that $(e,\mathbf{d})=m^{\prime}(e^{\prime},l\boldsymbol{\gamma})$ for some $m^{\prime}>1$. Then, (3.5.1) is an integer for some $l<m$, which implies $\mathbf{q}^{l\boldsymbol{\gamma}}=1$. ∎
@@
Let $e\in\mathbb{Z}$ and let $\mathbf{d}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$. Suppose that $\boldsymbol{\alpha}^{\prime}$ and $\boldsymbol{\alpha}\in I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$ are such that the interval
@@
does not meet any wall in $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$, and that $\boldsymbol{\alpha}$ is almost generic. Then, the following assertions hold:
@@
  • $\deg_{\boldsymbol{\alpha}}\mathscr{F}=0$, in which case $(\deg F,\mathbf{d}^{F})$ is a scalar multiple of $(e,\mathbf{d})$. (We regard $\mathbf{d}^{F}$ as an element of $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ by defining $d^{F}_{[j,i]}=\dim E_{[j,i]}\cap F_{p_{j}}$ for any $j,i]\in\boldsymbol{\tau}(\underline{\nu})$.)
@@
does not meet any wall in $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{d})$. Then, the morphism $\mathbf{Var}_{2}$ is well-defined and induces a bijection of $\mathbb{C}$-points.
@@
In case ($\mathbf{Aff}^{\infty}$), we can choose $\boldsymbol{\alpha}$ to be generic, and thus Proposition 3.7 applies, giving a well-defined $\mathbf{Var}_{2}$. A simplification in this case is that $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ consists of stable parabolic Higgs bundles, although $\boldsymbol{\alpha}^{\prime}$ is not necessarily generic. The second part of Proposition 3.7 shows that $\mathbf{Var}_{2}$ is bijective, since every object is stable in these moduli spaces. ∎
@@ 3.6. Forgetting filtered structures
$\mathbf{NH}_{2}:\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})\operatorname{\!\!\smash{\begin{array}[]{c}\sim\\[-10.00002pt] \longrightarrow\end{array}}\!\!}\mathcal{M}_{Dol}(\mathbf{d},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j}).$
Suppose that $\boldsymbol{\alpha}$ is an almost generic parabolic weight. Then, the eigenvalues $\boldsymbol{\xi}^{\prime}$ are almost generic, and thus the conjugacy classes $\mathcal{C}^{\prime}$ are almost generic. Similarly, if $\boldsymbol{\alpha}$ is generic, then the eigenvalues $\boldsymbol{\xi}^{\prime}$, and thus $\mathcal{C}^{\prime}$, are generic.
@@
Let $\mathbf{q}^{\prime}$ be the deformation parameter defined by $\boldsymbol{\xi}^{\prime}$ as in §1.1 and let $\boldsymbol{\gamma}\leq\mathbf{d}$ be a dimension vector. Suppose that $\mathbf{q}^{\prime\boldsymbol{\gamma}}=1$. We need to show that $\mathbf{d}$ is a multiple of $\boldsymbol{\gamma}$. Since the argument of $\xi_{[j,i]}$ is equal to $2\pi i\alpha_{[j,i]}$, the equality $\mathbf{q}^{\prime\boldsymbol{\gamma}}=1$ implies that
$\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}\gamma^{\ast}_{[j,i]}=-e^{\prime}$
for some $e^{\prime}\in\mathbb{Z}$, where we regard $\boldsymbol{\gamma}$ as an element of $\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ with $\gamma_{[j,0]}=\gamma_{\star}$ for any $j$. However, the weight $\boldsymbol{\alpha}$ is almost generic, and thus we have $m^{\prime}(e^{\prime},\boldsymbol{\gamma}^{\ast})=(e,\mathbf{d}^{\ast})$ for some $m^{\prime}\in\mathbb{Z}_{>0}$; in particular, $m^{\prime}\boldsymbol{\gamma}=\mathbf{d}$. The statement for generic $\boldsymbol{\alpha}$ is similar. ∎
@@
Let us show the surjectivity. Let $\rho_{1}\oplus\cdots\oplus\rho_{r}$ be a direct sum of simple representations $\rho_{i}$. Since $\mathbf{q}^{\prime}$ is almost generic, the dimension vector $\mathbf{d}^{(s)}$ of each $\rho_{s}$ is a multiple of $\mathbf{d}_{0}$, and thus $\mathbf{d}^{(s)}\cdot\boldsymbol{\theta}=0$. It follows that $\rho_{1}\oplus\cdots\oplus\rho_{r}$ is a $\boldsymbol{\theta}$-polystable representations, whence surjectivity.
Next, we show the injectivity. Let $\rho_{1}\oplus\cdots\oplus\rho_{r}$ be as above, and suppose that $\rho^{\prime}_{1}\oplus\cdots\oplus\rho^{\prime}_{t}$ is a direct sum of $\boldsymbol{\theta}$-stable representations $\rho^{\prime}_{s}$, which has $\rho_{1}\oplus\cdots\oplus\rho_{r}$ as its semi-simplification. Obviously, we have $t\leq r$. Let us show that $t<r$ is not possible. The dimension vector of each $\rho^{\prime}_{t}$ is also a multiple of $\mathbf{d}_{0}$. By replacing $\mathbf{d}$ by a smaller dimension vector if necessary, we may assume $t=1$; that is, the $\boldsymbol{\theta}$-stable representation $\rho^{\prime}_{1}$ has $\rho_{1}\oplus\cdots\oplus\rho_{r}$ as its semi-simplification. However, we have $\mathbf{d}^{(s)}\cdot\boldsymbol{\theta}=0$ for any $s$, contradicting the stability of $\rho^{\prime}_{1}$. We have shown that $t=r$. Then, each $\rho^{\prime}_{s}$ is necessarily simple, and thus is isomorphic to some $\rho_{s^{\prime}}$. The injectivity follows. We have also proved in the process that a $\boldsymbol{\theta}$-stable representation is necessarily irreducible, which proves the second statement of the proposition. ∎

### v2

@@ 3.5. Variation of parabolic weights
$\mathbf{Var}_{2}:\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})\longrightarrow\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j}).$
The types of $\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ and $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ are $\boldsymbol{\tau}(\underline{\nu})$ and $\boldsymbol{\tau}(\underline{\mu})$ respectively, with each $\mu_{j}\leq\nu_{j}$. We require that the following conditions are satisfied:
  • There is a degeneration of types $\sigma:\boldsymbol{\tau}(\underline{\mu})\longrightarrow\boldsymbol{\tau}(\underline{\nu})$ such that $\mathbf{d}^{\prime}=\sigma^{\ast}\tilde{\mathbf{d}}$.
  • The graded residue $\mathcal{O}_{j}$ is central in $\bigoplus_{i=0}^{\nu_{j}}\operatorname{\mathfrak{gl}}_{d^{\ast}_{[j,i]}}$ with $\nu_{j}+1$ distinct eigenvalues. Since the flag $E_{[j,\bullet]}$ at $p_{j}$ has length $\nu_{j}+1$, it consists precisely of sums of eigenspaces of $\Phi_{j}$.
@@
Remark 3.4.
(1). The filtered local systems in $\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi})$ also have type $\boldsymbol{\tau}(\underline{\nu})$. (2). The vectors $\mathbf{d}^{\ast}$ and $\tilde{\mathbf{d}}^{\ast}$ are equal up to permuting their components with the same $j$.
For an arbitrary parabolic weight $\boldsymbol{\alpha}$, a degeneration of parabolic structure may not preserve semi-stability. The usual option is to take a generic $\boldsymbol{\alpha}$ in a small neighbourhood of $\boldsymbol{\alpha}^{\prime}$; however, generic weights do not exist if the dimension vector is divisible.
@@
Fix $e\in\mathbb{Z}$ and strict $\mathbf{c}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ of rank $n$, and let $\boldsymbol{\alpha}\in\mathbb{R}^{\boldsymbol{\tau}(\underline{\nu})}$ be a parabolic weight such that
$e+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}c^{\ast}_{[j,i]}=0.$
We say that $\boldsymbol{\alpha}$ is almost generic if for any strict $\mathbf{c}^{\prime}\in\mathbb{Z}_{\geq 0}^{\boldsymbol{\tau}(\underline{\nu})}$ of rank $n^{\prime}\leq n$ and any $e^{\prime}\in\mathbb{Z}$ the equality
$e^{\prime}+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}c^{\prime\ast}_{[j,i]}=0$
holds if and only if the vector $(e,(c_{[j,i]})_{[j,i]})$ is a $\mathbb{Q}$-multiple of $(e^{\prime},(c^{\prime}_{[j,i]})_{[j,i]})$ (if and only if $(e,(c^{\ast}_{[j,i]})_{[j,i]})$ is a $\mathbb{Q}$-multiple of $(e^{\prime},(c^{\prime\ast}_{[j,i]})_{[j,i]})$); we say that $\boldsymbol{\alpha}$ is generic if the above equality does not hold unless $(e^{\prime},(c^{\prime}_{[j,i]})_{[j,i]})$ is zero or is equal to $(e,(c_{[j,i]})_{[j,i]})$.
@@
Note that we allow $\alpha_{[j,i]}=\alpha_{[j,i+1]}$ in this space. For $(e,\mathbf{c})$ as in the above definition, define
$I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{c}):=\{\boldsymbol{\alpha}\in I^{\boldsymbol{\tau}(\underline{\nu})}_{\Delta}\mid e+\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}\alpha_{[j,i]}c^{\ast}_{[j,i]}=0\},$
which is a hyperplane in $I^{\boldsymbol{\tau}(\underline{\nu})}_{\Delta}$. If a vector $(e^{\prime},\mathbf{c}^{\prime})$ is not a scalar multiple of $(e,\mathbf{c})$, then $I^{\boldsymbol{\tau}(\underline{\nu})}(e^{\prime},\mathbf{c}^{\prime})$ either does not meet $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{c})$ or intersects with $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{c})$ in a lower dimensional affine space, thus forming a wall in $I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{c})$. By definition, a weight $\boldsymbol{\alpha}\in I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{c})$ is almost generic if it lies in the complement of the union of these walls.
@@
$\sigma_{\ast}:I^{\boldsymbol{\tau}(\underline{\mu})}(e,\sigma^{\ast}\mathbf{c})\longrightarrow I^{\boldsymbol{\tau}(\underline{\nu})}(e,\mathbf{c}).$
Suppose that $\mathscr{E}=(E,E_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\nu})$. Then, a parabolic Higgs bundle $\mathscr{E}^{\prime}=(E,E^{\prime}_{[\bullet,\bullet]},\Phi)$ of type $\boldsymbol{\tau}(\underline{\mu})$ is obtained from the former by the degeneration $\sigma$ if $E^{\prime}_{[j,i]}=E_{[j,\sigma(i)]}$ for any $[j,i]$.
@@
Let $\mathscr{E}$ and $\mathscr{E}^{\prime}$ be as above, and let $\boldsymbol{\alpha}^{\prime}\in I_{\Delta}^{\boldsymbol{\tau}(\underline{\mu})}$. Then, we have
@@
Suppose that $(E,E_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\nu})$ and that $(E,E^{\prime}_{[\bullet,\bullet]},\Phi)$ is a parabolic Higgs bundle of type $\boldsymbol{\tau}(\underline{\mu})$ obtained from the former by the degeneration $\sigma$. Let $\boldsymbol{\alpha}^{\prime}$ be a parabolic weight for $E^{\prime}_{[\bullet,\bullet]}$. Let $F\subset E$ be a $\Phi$-invariant vector bundle, and let $F^{\prime}_{[\bullet,\bullet]}$ and $F_{[\bullet,\bullet]}$ be the parabolic structures induced by $E^{\prime}_{[\bullet,\bullet]}$ and $E_{[\bullet,\bullet]}$ respectively. Then,
@@
    $(F,F^{\prime}_{[\bullet,\bullet]},\Phi)$ is obtained from $(F,F_{[\bullet,\bullet]},\Phi)$ by a degeneration of types.
@@
Now, let $\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j})$ and $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ be as at the beginning of this subsection, and let $\sigma:\boldsymbol{\tau}(\underline{\mu})\longrightarrow\boldsymbol{\tau}(\underline{\nu})$ be the given degeneration. Let $e$ be determined by $\boldsymbol{\alpha}^{\prime}$ and $\mathbf{d}^{\prime}$ as in Lemma 3.2.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
In case ($\mathbf{Aff}^{\infty}$), the vector $\mathbf{d}$ is divisible, and so is $\tilde{\mathbf{d}}$ by Remark 3.4 (2). We can therefore choose $\boldsymbol{\alpha}$ to be generic, and thus Proposition 3.8 applies, giving a well-defined $\mathbf{Var}_{2}$. A simplification in this case is that $\mathcal{M}_{Dol}(\mathbf{d}^{\prime},\boldsymbol{\alpha}^{\prime},(\mathcal{O}^{\prime}_{j})_{j})$ consists of stable parabolic Higgs bundles (because the stability condition on the Betti side is chosen to be generic), although $\boldsymbol{\alpha}^{\prime}$ is not necessarily generic. The second part of Proposition 3.8 shows that $\mathbf{Var}_{2}$ is bijective, since every object is stable in these moduli spaces. ∎
@@ 3.6. Forgetting filtered structures
$\mathbf{NH}_{2}:\mathcal{M}_{B}(\mathbf{d},\boldsymbol{\beta},\boldsymbol{\xi}^{\prime})\operatorname{\!\!\smash{\begin{array}[]{c}\sim\\[-10.00002pt] \longrightarrow\end{array}}\!\!}\mathcal{M}_{Dol}(\tilde{\mathbf{d}},\boldsymbol{\alpha},(\mathcal{O}_{j})_{j}).$
Suppose that $\boldsymbol{\alpha}$ is an almost generic parabolic weight. Then, the eigenvalues $\boldsymbol{\xi}^{\prime}$ are almost generic, and thus the conjugacy classes $\mathcal{C}^{\prime}$ are almost generic. In particular, if $\boldsymbol{\alpha}$ is generic, then the eigenvalues $\boldsymbol{\xi}^{\prime}$ and thus $\mathcal{C}^{\prime}$ are generic.
@@
Let $\mathbf{q}^{\prime}$ be the deformation parameter defined by $\boldsymbol{\xi}^{\prime}$ as in §1.1 and let $\boldsymbol{\gamma}\leq\mathbf{d}$ be a strict dimension vector. Suppose that $\mathbf{q}^{\prime\boldsymbol{\gamma}}=1$. We need to show that $\mathbf{d}$ is a multiple of $\boldsymbol{\gamma}$. Let the argument of $\xi_{[j,i]}$ be $2\pi ia_{[j,i]}$, the equality $\mathbf{q}^{\prime\boldsymbol{\gamma}}=1$ implies that
$\sum_{[j,i]\in\boldsymbol{\tau}(\underline{\nu})}a_{[j,i]}\gamma^{\ast}_{[j,i]}=-e^{\prime}$
[…]
@@
Let us show the surjectivity. Let $\rho_{1}\oplus\cdots\oplus\rho_{r}$ be a direct sum of simple representations $\rho_{s}$. Since $\mathbf{q}^{\prime}$ is almost generic, the dimension vector $\mathbf{d}^{(s)}$ of each $\rho_{s}$ is a multiple of $\mathbf{d}_{0}$, and thus $\mathbf{d}^{(s)}\cdot\boldsymbol{\theta}=0$. It follows that $\rho_{1}\oplus\cdots\oplus\rho_{r}$ is a $\boldsymbol{\theta}$-polystable representation, whence surjectivity.
Next, we show the injectivity. Let $\rho_{1}\oplus\cdots\oplus\rho_{r}$ be as above, and suppose that $\rho^{\prime}_{1}\oplus\cdots\oplus\rho^{\prime}_{t}$ is a direct sum of $\boldsymbol{\theta}$-stable representations $\rho^{\prime}_{s}$, which has $\rho_{1}\oplus\cdots\oplus\rho_{r}$ as its semi-simplification. Obviously, we have $t\leq r$. Let us show that $t<r$ is not possible. The dimension vector of each $\rho^{\prime}_{s}$ is also a multiple of $\mathbf{d}_{0}$. By replacing $\mathbf{d}$ by a smaller dimension vector if necessary, we may assume $t=1$; that is, the $\boldsymbol{\theta}$-stable representation $\rho^{\prime}_{1}$ has $\rho_{1}\oplus\cdots\oplus\rho_{r}$ as its semi-simplification. However, we have $\mathbf{d}^{(s)}\cdot\boldsymbol{\theta}=0$ for any $s$, contradicting the stability of $\rho^{\prime}_{1}$. We have shown that $t=r$. Then, each $\rho^{\prime}_{s}$ is necessarily simple, and thus is isomorphic to some $\rho_{s^{\prime}}$. The injectivity follows. We have also proved in the process that a $\boldsymbol{\theta}$-stable representation is necessarily simple, which proves the second statement of the proposition. ∎